OpenStudy (anonymous):

Katrina is fencing her yard, which is shaped as a rhombus. She used (5x + 1) meters of fencing material for one side and (7x - 11) meters for the adjacent side. What is the perimeter of the yard?

OpenStudy (anonymous):

We know Rhombus = all sides equal. So, equate the two sides, find x. Then, find value of one of the sides. Multiply by 4 for perimeter.

OpenStudy (aroub):

The perimter of a rhombus is 4 times the length of any one side. & the sides of a rhombus are always congruent. 5x+1=7x-11 -7x+5x=-11-1 -2x=-12 x=6 5(6)+1 31 The perimeter: 4(31)=124

OpenStudy (anonymous):

Thanks guys! can i get help on this one...